OpenStudy (kendricklamar2014):

A radio signal travels at 3.00 * 10^8 meters per second. How many seconds will it take for a radio signal to travel from a satellite to the surface of Earth if the satellite is orbiting at a height of 3.54 * 10^7 meters? Show your work.

hero (hero):

Divide Remember \[Time = \frac{Distance}{Rate}\]

OpenStudy (kendricklamar2014):

So 3.00/10^8

hero (hero):

\[T = \frac{3.54 \times 10^7}{3.00 \times 10^8}\] \(= \dfrac{3.54}{3.00} \times \dfrac{10^7}{10^8}\)

OpenStudy (kendricklamar2014):

So u times 3.54 x 3.00?

hero (hero):

Does \(\dfrac{3.54}{3.00}\) look like multiplication to you?

OpenStudy (kendricklamar2014):

sorry i meant the answer of 3.54/3.00 x the answer of 10^7/10^8

OpenStudy (kendricklamar2014):

3.54/3.00= 1.18, correct?

hero (hero):

Now we're getting somewhere.

OpenStudy (kendricklamar2014):

ok so 10^7/10^8= 1^-1 or 1^1?

hero (hero):

Not exactly

hero (hero):

The base remains the same.

hero (hero):

\(a^b/a^c = a^{b - c}\) You only subtract the exponents.

OpenStudy (kendricklamar2014):

So 10^-1?

OpenStudy (kendricklamar2014):

@Hero

hero (hero):

Yes, so your result is 1.18 x 10^{-1} but don't forget units.

OpenStudy (kendricklamar2014):

So the answer would be 10.18^-1

hero (hero):

No. I just wrote what the final result is. You're changing it to something else for no reason.

OpenStudy (kendricklamar2014):

So 1.18 x 10^{-1} is the answer?

OpenStudy (kendricklamar2014):

@Hero

hero (hero):

The time is measured in seconds, so the final result should be expressed as 1.18 × 10^{-1} seconds.

OpenStudy (kendricklamar2014):

thank u so much :D